OpenStudy (anonymous):

hey, im in 9th and need help with algebra; trinomials. if you cold help me that would be great. thanks,

OpenStudy (anonymous):

What do you need help with, specifically?

OpenStudy (anonymous):

r(square) + 2r = 24

OpenStudy (anonymous):

Solving for r? There are so many ways of doing this it's not even funny. Here are a few: Quadratic equation: http://en.wikipedia.org/wiki/Quadratic_equation r^2 + 2r - 24 = 0 so r = [-2 ± sqrt(2^2 - 4*1*(-24))] / [2*1] r = 4 or -6 Completing the square: r^2 + 2r = -24 r^2 + 2r + (2/2)^2 = 24 + (2/2)^2 r^2 + 2r + 1 = 25 (r + 1) ^ 2 = 25 r + 1 = ±sqrt(25) r + 1 = ±5 r = -1 ±5 r = 4 or -6 Seat-of-the-Pants Factoring: r^2 + 2r - 24 = 0 (try to find two numbers that sum to 2 and multiply to -24) -1 * 24 = -24 but they don't sum to 2 -3 * 8 = - 24 but they dont' sum to 2 -4 * 6 = - 24 and they do sum to 2. That's the set we need. so this expression can be factored like so: (r^2 + 2r - 24 = (r - 4)(r+6) And this new expression will only equal zero if one of the terms in parentheses equals zero. That'll only happen at r = 4 or -6